OpenStudy (anonymous):

PLEASE HELP ME! 3^-3x=17 sole for x using the natural log function

OpenStudy (anonymous):

\[-3x=\frac{\ln(17)}{\ln(3)}\] is a start

OpenStudy (anonymous):

then \[x=-\frac{\ln(17)}{3\ln(3)}\]

OpenStudy (anonymous):

don't never forget, solve \[b^x=A\] via \[x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

what happend to the -3?

OpenStudy (anonymous):

in the 2nd step

OpenStudy (anonymous):

i divided by it in the second step

OpenStudy (anonymous):

i divided by 3, put the minus sign out front

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

so what does x equal?